\int \:2^{2x}\cdot 5^xdx
1
Expert's answer
2020-09-14T19:19:24-0400
"\int2^{2x}\cdot5^xdx=\int4^x\cdot5^xdx=\int20^xdx=\dfrac{20^x}{\ln(20)}+C"